OpenStudy (konradzuse):

Find the area of the region bounded by the graphs of the equation y = x^3 + x [4,0]

OpenStudy (konradzuse):

I don't think I do an integration? \[\int\limits_{0}^{4} x^3 +x?\]

OpenStudy (anonymous):

we have to do integration here

OpenStudy (konradzuse):

oh okay was my setu correct?

OpenStudy (anonymous):

yes correct!!:)

OpenStudy (anonymous):

do we have to find the area between x=0 to x=4?

OpenStudy (konradzuse):

so x^4/4 + x^2/2

OpenStudy (konradzuse):

x = 4 y =0

OpenStudy (konradzuse):

Find the area of the region bounded by the graphs.

OpenStudy (anonymous):

\[\int\limits\limits_{0}^{4} x^3 +x\] is the area under x=0 and x=4

OpenStudy (konradzuse):

ok so what do I do for x = 4 and y = 0?

OpenStudy (agent47):

ok... soo

OpenStudy (konradzuse):

:)

OpenStudy (agent47):

what do you mean by [4,0]?

OpenStudy (agent47):

HElloooooo are you there?

OpenStudy (konradzuse):

I'm sorry.

OpenStudy (konradzuse):

[4,0] meaniong x =4 y = 0.

OpenStudy (agent47):

ohh ok

OpenStudy (agent47):

So just integrate that equation from 0 to 4.

OpenStudy (konradzuse):

So like I posted from the start? lawls...

OpenStudy (agent47):

Int(x^3+x,x,0,4)=x^4/4+x^2/2=4^3+8

OpenStudy (konradzuse):

\[\int\limits_{0}{4} \frac{x^4}{4} \frac{x^2}{2}\]

OpenStudy (agent47):

NOOOOOOOOOOOOOOOOOO

OpenStudy (konradzuse):

\frac{x^4}{4} \frac{x^2}{2}\[ \frac{4^4}{4} \frac{4^2}{2} = 4^ \frac{^2}{2}\]

OpenStudy (konradzuse):

nooo? :(

OpenStudy (agent47):